4. Suppose you have the following estimation results from the monthly data of January 1960 to December 1972: yt = 2.03 +0.152where the standard errors are in parentheses, Residual Sum of Squares (RSS)=18.48, and Explained Sum of Squares (ESS)=109.60.
Answer #1

au Ytt 2.034 0.152 at (12] (0.088) @ Sample size 13 6 R2 a Ess - where fss = - Explained sum of equale see TIS = Total sum of+ - 203+0 • IS 2224 -1-2}___(00038) Hor- 250. HAB <0.|| since D is small, 10=13 By t - Statistica ____4 = [-EB) कि - 0.152-Since value lies in Rejection region Hence well trypothesis rejected Hence B coil o quy. Confidence Interval will be given byPogo o 95 for 192 95% confidence griferral — -64. №10 6-е іn h- Tule have _= 0.152 then you a Thovefore - 2B + 1 = 2(0.152) E